OpenStudy (anonymous):

How do i answer this What percent of 125 is 75?

OpenStudy (ipwnbunnies):

\[\frac{ part }{ whole } * 100%\]

OpenStudy (anonymous):

\[75/125 = x/100\] simple algebra

OpenStudy (anonymous):

so 75 times 100 divided by 125 = your answer

OpenStudy (anonymous):

Thanx guys

OpenStudy (anonymous):

No problem